OpenStudy (johnjules2000):

Last question: 2a -7 = -9 I have to solve for a.

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

\[2a-7=-9\]\[2a=-2\]\[a=-1\]

OpenStudy (anonymous):

2a−7=−9 2a=−2 a=−1

OpenStudy (johnjules2000):

I did this: 2a - 7 = -9 \ add 7 to both sides 2 a = -2 \ multiply by 2 What am I doing wrong??

OpenStudy (anonymous):

you dont multiply 2 you divide by 2

OpenStudy (anonymous):

so it is -2/2 which is -1

OpenStudy (johnjules2000):

Ok, thank you. I am trying to explain to my daughter & couldn't make it work. Thank you!

OpenStudy (anonymous):

oh ok no problem!!!